OpenStudy (anonymous):

A line has a slope of 1 2 and passes through the point (0,5). What is the equation of the line?

OpenStudy (anonymous):

Use the point slope form of a line \[(y-y _{1})=m(x-x _{1})\]

OpenStudy (anonymous):

has a slope of * 1/2

OpenStudy (anonymous):

where \[y _{1}=5 and x _{1}=0 and m=1/2\]

OpenStudy (anonymous):

simpify: (m + 11) + (n + 44)

OpenStudy (anonymous):

A)m + n + 84 B)m + n + 55 C)mn + 55 D)mn + 484

OpenStudy (anonymous):

B

OpenStudy (anonymous):

i love you

OpenStudy (anonymous):

lol did you get the first problem?

OpenStudy (anonymous):

Which expression is equivalent to x3 – x2 – 14x + 24? A)(x – 2)(x – 3)(x + 4) B)(x – 2)(x + 3)(x – 4) C)(x + 2)(x – 3)(x – 4) D)(x + 2)(x + 3)(x + 4)

OpenStudy (anonymous):

yes, thank u

OpenStudy (anonymous):

are you sure the problem is not (x-3)-(x-2)-(14x+24) ??

OpenStudy (anonymous):

yeah its that

OpenStudy (anonymous):

wait it's \[x ^{3}-x ^{2}-14x-32\] right???

OpenStudy (anonymous):

x = -2y + 300 3y = 800 - 4x Which statement about the pair of equations is true? A)The equation pair has no solutions. B)The equation pair has many solutions. C)The equation pair has exactly 1 solution. D)The equation pair has exactly 2 solutions.

OpenStudy (anonymous):

oppps +24

OpenStudy (anonymous):

yeah yeah thats a selection

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

d

OpenStudy (anonymous):

A bike rental company charges $50 for the first hour, and $25 for each additional hour. Michael must spend no more than $200 on his bike rental for the day. What is the maximum number of hours that Michael can rent a bike? A)3 hours B)4 hours C)5 hours D)7 hours